OpenStudy (anonymous):

If xis a number in the interval (4,8) state all intergers that satisfy the given quantity 7x-3(4x-8) ≤ 6x +12 -9x

OpenStudy (er.mohd.amir):

7x-12x+24<=-3x+12 -5x+3x<=12-24 -2x<=-12 x>=6 that is x=7

OpenStudy (anonymous):

good morning

OpenStudy (anonymous):

^^ Er.Mohd.AMIR helped you out